Minute of KAGRA commissioning meeting
- Date: 2018/4/3 11AM~12PM
- Attendance: Uchiyama, Kokeyama, Yokozawa, Aso, Izumi, Nakano
- eZuce: Miyakawa, Yamamoto(Kazuhiro), Suzuki, Hirose, Yamamoto, Inoue, Ushiba, Shoda, Kimura, Miyoki, Tomaru
1. Report of work of the last week and plan of work
- IFO: Izumi
- Beam from both X and Y end was confirmed at REFL port.
- ETMs: Ushiba
- ETMY damping mirror motion has been suppressed to ~1urad in yaw (Miyamoto).
- ETMX has no heat link and is characterized without heat link. Heat link is connected on Thursday.
- RT model
- Cable connection is almost done.
- Dolphin should be connected as soon as possible. It can stop all models connected Delphin.
- BS model will be splitted and one of them will be down sampled to 4kHz.
- Oplev: Ushiba
- Test mas calibration was done.
- Length oplev is not ready yet.
- Marionette Oplev will be calibrated when door closed.
- CAL: Yamamoto, Yokozawa
- RT calibration using IMC as training of calibration will be tested soon.
- Off line calibration using FIR filter won't be used in phase 1.
- k1tw0 is sometimes down as like other FWs.
- SSD will be tested soon.
2. IFO check list
- It will be prepared soon.
- Please reply accommodation mail from Uchiyama-san.
- Think for emergency.
4. Tasks or discussions
DAQ fix -> Miyakawa, Yamamoto
Selection of full channel -> DGS Yamamoto
Selection of science channel -> Commissioning Izumi
- Screens at control room
- Needs laser safety information.
GW alert: online alert in the control room might be prepared. -> Togoshi
- Dust monitor using sapphire disk will be implemented around 4/10.
- Phase 1 summarize should talk wide range view.
- Type A has double stable point.
- Monitor for water level, and vacuum gage for weekend should be implemented.
- Patrol light should be implemented soon, even manual one.
- Gardian of TYPE A must be implemented and tested.
- Pump down meeting meeting at Fan hut: 12:30 at 4/5, 12:30 at 4/11
- Next meeting: 4/10